# A Quality Diversity Method to Automatically Generate Multi-Agent Path Finding Benchmark Maps

### v4

@@ 3.1 Objective
In all experiments, our objective is a function $f:\boldsymbol{X}\rightarrow\mathbb{R}$, where $\boldsymbol{X}$ is the space of all maps. The function $f$ runs one or two MAPF algorithms on $N_{e}$ different MAPF instances with $N_{a}$ agents. For CBS, EECBS, and PBS, we set a time limit $T$ for each run. For PIBT and LTF, we set a maximum makespan $M$.
@@
One-Algorithm Experiments. We intend to generate maps that are challenging for the MAPF algorithms in one-algorithm experiments. The objective quantifies the empirical hardness of a map $\boldsymbol{x}\in\boldsymbol{X}$ to a MAPF algorithm $\phi\in A$, where $A=\{\text{CBS},\ \text{EECBS},\ \text{PBS},\ \text{PIBT},\ \text{LTF}\}$ is all MAPF algorithms considered in the experiments.
@@
For CBS, EECBS and PBS, we maximize their CPU runtime because they are exponential-time algorithms with the major limitation being frequently running out of time. We do not consider the solution quality as the objective because CBS and EECBS are proven to be optimal and bounded-suboptimal algorithms. PBS, while being unbounded suboptimal, empirically finds close-to-optimal solutions. For example, Ma et al. (2019) shows that PBS finds solutions never over 4% worse than optimal in fixed game maps. Therefore, we maximize the CPU runtime to find failure cases for CBS, EECBS, and PBS. Concretely, suppose the function $t_{\phi}:\boldsymbol{X}\rightarrow\mathbb{R}_{>0}$ computes the average CPU runtime by running $\phi\in A$ in $N_{e}$ instances, the objective $f$ for the one-algorithm experiments is $f(\boldsymbol{x})=t_{\phi}(\boldsymbol{x}),\phi\in\{\text{CBS},\ \text{EECBS},\ \text{PBS}\}$. If the algorithms fail to return solutions within the timelimit $T$, we set the CPU runtime as $T$.
@@
$SR_{\phi}:\boldsymbol{X}\rightarrow\mathbb{R}_{>0}$ computes the success rate of running algorithm $\phi$ on one MAPF instance of a map, which is computed as the percentage of agents that successfully reach their goals within the given makespan $M$. $SoC_{\phi}:\boldsymbol{X}\rightarrow\mathbb{R}_{>0}$ computes the sum-of-cost of the solution if $SR=1$. $C$ is a large constant making sure that the RSR score of maps with $SR=1$ always dominates that of maps with $SR<1$. Intuitively, if not all agents reach their goals ($SR<1$), we maximize the success rate. If all agents reach their goals ($SR=1$), we regularize the success rate with the sum-of-cost to quantify the quality of the solution. Let the function $r_{\phi}:\boldsymbol{X}\rightarrow\mathbb{R}_{>0}$ return the average regularized success rate by running algorithm $\phi$ in $N_{e}$ MAPF instances. Concretely, $r_{\phi}(\boldsymbol{x})=\sum_{i=1}^{N_{e}}RSR_{\phi}^{(i)}(\boldsymbol{x})$, where $i$ denotes the $i$-th run of algorithm $\phi$. Then the objective of PIBT and LTF are $f(\boldsymbol{x})=r_{\text{PIBT}}(\boldsymbol{x})$ and $f(\boldsymbol{x})=r_{\text{LTF}}(\boldsymbol{x})$, respectively. We do not consider runtime as the objective of PIBT and LTF because they run very fast, making CPU runtime not a good metric of their performance in a map.
@@
Two-Algorithm Experiments. To compare two given MAPF algorithms in two-algorithm experiments, we aim to generate maps that maximize the performance gap between the two algorithms. We consider two pairs of comparisons: (1) EECBS vs. PBS, and (2) PIBT vs. LTF. We compare EECBS and PBS because (1) both suboptimal algorithms use a two-level search based on collision avoidance, (2) both find near-optimal solutions empirically, and (3) few prior works have compared them thoroughly. We compare PIBT and LTF because (1) both use a pre-defined policy, either rule-based or learned, to move the agents step by step to their goals while avoiding collisions, (2) both run fast but suffer from deadlocks, and (3) while the LTF work Skrynnik et al. (2024) has compared them on warehouse maps, showing that LTF outperforms PIBT, we want to find cases where PIBT outperforms LTF.
To compare EECBS and PBS, we set the objective as the absolute difference in the average CPU runtime. Concretely, the objective of a generated map $\boldsymbol{x}\in\boldsymbol{X}$ is computed as $f(\boldsymbol{x})=|t_{\text{EECBS}}(\boldsymbol{x})-t_{\text{PBS}}(\boldsymbol{x})|$. Similarly, to compare PIBT and LTF, we set the objective as the absolute difference in the regularized success rate, i.e., $f(\boldsymbol{x})=|r_{\text{PIBT}}(\boldsymbol{x})-r_{\text{LTF}}(\boldsymbol{x})|$.
3.2 Diversity Measures
@@
We aim to select diversity measures such that the generated maps are diverse in terms of (1) general hardness towards all algorithms to analyze whether the general hardness of a map aligns with the hardness to a specific algorithm, and (2) spatial arrangement of the obstacles to generate maps of different patterns.
@@
General Hardness. To measure the general hardness of a map, the most commonly used metric is the number of obstacles. The number of obstacles in a map affects the space available to resolve collisions as well as the search space of the algorithms. Therefore, diversifying it would give a range of difficulties. We also consider two metrics that are proposed specifically to reflect the hardness of a map: the standard deviation of Betweenness Centrality (std of BC) Ewing et al. (2022) and $\lambda_{2}$ Ren et al. (2024). To compute the std of BC, we search for the shortest path between every pair of empty spaces in the map and then calculate the standard deviation of the usage of all empty spaces. The higher the std of BC, the more congestion and the harder the MAPF problem. To compute $\lambda_{2}$ of a map, we view the map as a graph $G(V,E)$, and compute the second-smallest eigenvalue of the normalized Laplacian matrix of $G$. Ren et al. (2024) shows empirically that $\lambda_{2}$ is correlated with the hardness of the map.
@@
Spatial Arrangement. Aside from the hardness of the maps, we want to diversify the spatial arrangement of the maps, making the generated maps stylistically diverse. We first consider the KL divergence of the tile pattern distribution Fontaine et al. (2021b) between the tile pattern distribution of the generated map and one category of fixed maps selected from the MAPF benchmark Stern et al. (2019). A tile pattern is one possible arrangement of the obstacles and empty spaces in a 3 $\times$ 3 grid. To compute the tile pattern distribution of a map, we count the number of all possible tile patterns in the map. We then compute the KL divergence between the tile pattern distribution of the map and a fixed distribution computed from all the maze maps [^2] https://movingai.com/benchmarks/maze/index.html of the MAPF benchmark set Stern et al. (2019). The smaller the KL divergence of tile pattern distribution, the more similar in local patterns our generated map is to the selected set of benchmark maps. We choose maze maps because they empirically yield the most complex patterns. In addition, we consider the entropy of the tile pattern distribution Zhang et al. (2023a). The lower the entropy, the more regularized patterns the map possesses. Finally, we consider the KL divergence of the Weisfeiler-Lehman (WL) graph feature Shervashidze et al. (2011) between the generated map and the maze maps. Similar to the tile pattern distribution, a smaller KL divergence of the WL graph feature implies more similarities between the local patterns of the generated map and the maze maps.
@@
In our experiments, we choose the number of obstacles and KL divergence of tile pattern distribution as the diversity measures. We empirically observe that, among the measures considered, the number of obstacles is the most correlated with the general hardness of the maps, and the KL divergence of tile pattern distribution yields the most diverse map patterns. We justify our choice with experimental results in Appendix A.
4 Benchmarking One Algorithm
@@
For one-algorithm experiments, we intend to generate a diverse set of difficult benchmark maps targeted for CBS, EECBS, PBS, PIBT, and LTF.
@@
Figure 2: Examples of one-entry space and one-tile entry.
Figure 3: Archive for CBS with representative maps.
4.1 Experiment Setup
@@
Hyperparameters. Defining $w$ as the suboptimality bound away from the optimal solution, we use $w=1.5$ for EECBS. For CBS, we use the implementation of EECBS with $w=1$. We set $T=20$ seconds for CBS, EECBS, and PBS, $M=1000$ for PIBT, and $M=512$ for LTF. For all algorithms, we fix the map size as 32 $\times$ 32 and the number of agents at 150. The pre-defined range for the number of obstacles is $[O_{lb}=307,O_{ub}=717]$, which corresponds to 30% to 70% of the map size. For each evaluation, we run the given MAPF algorithm in $N_{e}=5$ different instances. For all algorithms, we evaluated $N_{eval}=10,000$ maps. For CBS, EECBS, and PBS, we run evaluations with $N_{a}=50$ agents, and for PIBT and LTF, we use $N_{a}=150$. We use more agents for PIBT and LTF because they empirically have higher success rates in congested MAPF instances. We include implementation details and compute resources in Appendix B.
4.2 Results
Generated Maps. We first define one-entry space and one-tile entry (shown in Figure 2) as features of the benchmark maps. A one-entry space represents the empty space surrounded by three obstacles. A one-tile entry represents a single empty space separating two corridors. In Figures 3, 4 and 5, the yellow and dark blue cells indicate maps with high and low CPU runtime, respectively. In Figures 6 and 7, the yellow and dark blue cells indicate maps with low and high success rates, respectively. The colored tiles in the maps display 10 out of 150 pairs of start and goal locations. While we are primarily interested in maps that are challenging for the algorithms, we also present easy maps as the by-products of the experiments. We summarize the typical patterns of the maps and show more maps with similar patterns in Appendix C.
Figure 4: Archive for EECBS with representative maps.
CBS. Figure 3 shows the archive of maps for CBS. It is easy to generate challenging maps for CBS. Map (a), Map (c), and Map (d) show the most typical hard maps for CBS. Map (a) and Map (c) contain many long corridors, and Map (d) contains many one-entry spaces with long corridors. Meanwhile, we also find that CBS can easily solve maps with fewer obstacles and a large chunk of empty spaces, such as Map (b). It provides more space for agents to avoid collisions.
EECBS. Figure 4 shows the generated maps for EECBS. Compared to CBS, it is harder to generate challenging maps for EECBS. Map (b) and Map (c) are typical hard maps for EECBS. Map (b) contains a large chunk of empty space, but with many long corridors and one-tile entries. Map (c) contains many long corridors and one-tile entries. EECBS runs out of time on both maps. On the other hand, EECBS performs well in maps with more empty spaces between each long obstacle component (Map (a)) and maps with short obstacle components (Map(d)), where obstacle components refer to connected components of two or more obstacles.
Figure 5: Archive for PBS with representative maps.
PBS. Figure 5 shows the archive of maps for PBS. We discover that PBS returns no solutions in many maps. Therefore, as a side discovery, we show the maps in which PBS returns no solution in at least one out of five MAPF instances during the evaluation in the archive as orange cells. Map (c) and Map (d) are typical hard maps for PBS. Map (c) contains many long corridors with one-tile entries in between, and Map (d) contains many one-entry spaces. PBS always reports no solution on Map (c) and runs out of time on Map (d). On the other hand, Map (a) and (b) show two typical easy maps for PBS. Map (a) contains long corridors but with more entry spaces in between. Map (b) contains many short obstacle components, providing more space for agents to avoid collisions.
Figure 6: Archive for PIBT with representative maps.
PIBT. Figure 6 shows the archive of maps of PIBT. The success rate of PIBT decreases with an increased number of obstacles and KL divergence. Similar to EECBS and CBS, PIBT performs poorly in maps with long corridors, such as Map (c). With a similar number of obstacles and smaller KL divergence, we can find easy maps for PIBT. They have large chunks of empty spaces with few corridors. Map (a) is one such example. In the middle of the archive, we observe maps with success rates of around 50%. They usually have long corridors and one-entry spaces. Map (d) is an example. In the bottom left corner, we find Map (b), which has many one-entry spaces but fewer obstacles, achieving a success rate of 100%. Overall, PIBT can efficiently solve maps that include long corridors and one-entry spaces if given sufficient space.
Figure 7: Archive for LTF with representative maps.
LTF. Figure 7 shows the archive of maps for LTF. Similar to PIBT, LTF performs worse with more obstacles and higher KL divergence, as shown in Map (a). Similar to challenging maps for CBS, EECBS, and PIBT, Map (a) contains long corridors. With fewer obstacles, LTF performs poorly on maps with many one-entry spaces, such as Map (b) and Map (d). These two maps contain different numbers of obstacles, but they have similar patterns of one-entry spaces. On the other hand, LTF performs better with fewer obstacles. These maps possess large chunks of empty spaces, such as Map (c).
Validation Results. We only evaluate each map by running $N_{e}=5$ MAPF instances during the QD search. To further validate our observations about the hardness of the maps, we additionally present validation results in Table 1 of Appendix D by running 200 MAPF instances in the selected maps. The results validate the hardness of the maps.
5 Comparing Two Algorithms
Given the different categories and properties of the MAPF algorithms, hard maps for one algorithm might not be hard for the other. Therefore, for two-algorithm experiments, we aim to generate maps that are easy for one algorithm and hard for the other by maximizing the performance gap between them.
Figure 8: Archive for EECBS vs PBS with representative maps.
Figure 9: Archive for PIBT vs LTF with representative maps.
5.1 Experiment Setup
We run EECBS with $w=1.5$. For EECBS and PBS, we use $T=20$ seconds. For PIBT and LTF, we use $M=512$, which is used to train the LTF policy Skrynnik et al. (2024). We use the same $O_{lb}$, $O_{ub}$, $N_{e}$, $N_{eval}$, $N_{a}$, and map sizes as the one-algorithm experiments.
5.2 Results
Generated Maps. We show the archives and representative benchmark maps in Figures 8 and 9. To better visualize the comparisons of the algorithms, we plot the differences, instead of the absolute differences, between the CPU runtime or success rate of the algorithms. In Figure 8, the yellow and dark blue cells represent EECBS having lower and higher CPU runtime than PBS, respectively. In Figure 9, the yellow and blue cells represent PIBT having higher and lower success rates than LTF, respectively.
EECBS vs. PBS. Figure 8 shows the archive comparing EECBS and PBS. In most cases, EECBS and PBS have similar performances. Maps with more than 512 obstacles are usually too hard for both algorithms. With fewer than 400 obstacles, however, EECBS generally performs better than PBS in two types of maps, exemplified in Map (a) and (d). Map (a) is similar to Map (a) of Figure 4, with long but wide corridors and one-entry spaces between adjacent corridors. Map (d) shows a random pattern with many one-entry spaces. On the other hand, Map (b) and (c) are examples where PBS outperforms EECBS. Notably, while Map (c) contains narrow long corridors, both ends of the corridors are open, potentially making PBS resolve conflicts easier.
Our results serve as a more comprehensive comparison between EECBS and PBS. Notably, only one work Chan et al. (2023) has systematically compared EECBS and PBS. They conclude that EECBS outperforms PBS in five out of six maps selected from the MAPF benchmark Stern et al. (2019). In contrast, our results indicate that EECBS and PBS exhibit similar behavior on most maps, with each having distinct advantages on different map patterns.
PIBT vs. LTF. Figure 9 shows the archive comparing PIBT and LTF. In most cases with less than 50% of obstacles, PIBT and LTF have a similar success rate. In maps with large chunks of empty space such as Map (d), both PIBT and LTF perform well. With more than 50% of obstacles, PIBT generally has a 5% to 10% higher success rate on average than LTF. In most cases, these maps contain many one-entry spaces, such as Map (a). On the other hand, the maps in which LTF outperforms PIBT are concentrated in the upper right and bottom left, such as Map (b) and (c).
Our results provide additional comparisons between PIBT and LTF to the experiments conducted by Skrynnik et al. (2024), which conclude that LTF outperforms PIBT by comparing them on a set of human-designed warehouse maps.
Validation Results. Similar to Section 4, we run 200 MAPF instances in selected maps and show the results in Table 2 of Appendix D to validate the hardness of the generated maps.
6 Discussion and Conclusion
We propose a framework based on QD algorithms to generate diverse benchmark maps for MAPF algorithms. We provide two concrete realizations of the framework, presenting experimental results with diverse benchmark maps of different levels of hardness and patterns. For the one-algorithm experiments, we generate hard benchmark maps for five representative MAPF algorithms by either maximizing the CPU runtime or minimizing the regularized success rate. For the two-algorithm experiments, we generate benchmark maps that are hard for one algorithm and easy for the other.
When and How to Use the Proposed Framework. While developing novel MAPF algorithms, researchers can use our one-algorithm pipeline to find maps their algorithm has difficulty solving. By generating such maps, they can have a better understanding of the weaknesses of their algorithms. Researchers can also leverage our two-algorithm pipeline to compare their proposed algorithm with an existing algorithm. They can observe which kinds of maps their algorithm performs better or worse than the existing algorithm. We argue that this is a more systematic way of comparing two algorithms than comparing on a fixed set of benchmark maps because researchers might intentionally or unintentionally cherry-pick maps that favor their algorithm or overfit the design of their algorithms to the selected set of maps.
Nevertheless, the pipelines shown in the paper are not the only realizations of our framework. Researchers can tailor the framework for alternative purposes by designing the objectives and diversity measures. For example, researchers can minimize, instead of maximize, the CPU runtime of CBS, EECBS, and PBS with a larger number of agents to generate maps that are easy, instead of hard, for these algorithms. While we present some easy maps for each selected algorithm in Section 4, they are by no means an exhaustive set. If easy maps are of interest, researchers should use alternative objectives, such as minimizing the CPU runtime or maximizing the success rate. If researchers are interested in both hard and easy maps, using the CPU runtime or regularized success rate as diversity measures is a better choice.
Researchers can also generalize the two-algorithm pipeline to an arbitrary number of MAPF algorithms to rank the performance of more than two algorithms. In addition, our framework focuses on generating MAPF benchmark maps. Future research can explore generating diverse MAPF instances by incorporating numbers of agents, map sizes, as well as start and goal locations into the pipeline.
Computational Cost. Our proposed framework relies on QD algorithms, which require a massive number of evaluations in the generated benchmark maps. Specifically, with our one-algorithm experiment setup, it takes 6 to 7 hours to generate benchmark maps for PIBT and LTF, 10 hours for PBS, and more than 12 hours for CBS and EECBS. Future works can focus on reducing the computational cost of QD algorithms by using surrogate models Zhang et al. (2022). Meanwhile, we will release our most representative generated benchmark maps online. If the researchers’ new algorithm is built on one of the algorithms we have tested, they can directly evaluate new algorithms on our benchmark maps.
Appendix A Choosing Diversity Measures
In this section, we justify our choice of diversity measures by setting up the archive with every two pairs of competitive diversity measures and exploring their relationship and behavior. We show the following experiments: (1) we compare the number of obstacles and standard deviation of Betweenness Centrality (std of BC)  Ewing et al. [2022], the number of obstacles and $\lambda_{2}$ Ren et al. [2024], and show that the number of obstacles could quantify the general hardness of the generated maps better than std of BC and $\lambda_{2}$. (2) We compare the KL divergence of the tile pattern distribution and map entropy, the KL divergence of the tile pattern distribution and the KL divergence of the WL graph feature, and show that the KL divergence of the tile pattern distribution can generate more complex patterns than map entropy and has less computational cost than the KL divergence of WL graph feature.
Figure 10: Scatter plot of about 4370 generated benchmark maps.
A.1 General Hardness
Number of Obstacles vs. Std of BC. To compare the ability to quantify the general hardness of the maps between the number of obstacles and the std of BC, we generate 4370 maps with our pipeline with the objective set as similarity score Zhang et al. [2023a] between repaired and unrepaired maps. We then set the diversity measure as the number of obstacles and std of BC. In this way, we generate a collection of benchmark maps that are not biased towards any algorithm while making sure that they are diverse in terms of the two measures of interest. Then we run PBS on all the generated maps, collecting average CPU runtime over $N_{e}=5$ MAPF instances and the std of BC. We plot them as a scatter plot in Figure 10. We observe that in the y-axis, with the increase of obstacles, the CPU runtime of PBS is increasing, which is consistent with the one-algorithm experiment result of PBS in Section 4 of the main text. With fewer variations in the CPU runtime of PBS at a given number of obstacles, we can quantify the general hardness of the maps based on the number of obstacles. From the x-axis, however, there are many maps with the same std of BC but showing different CPU runtimes of the maps for PBS. Therefore, we argue that the number of obstacles is a better measure than the std of BC to quantify the general hardness of the maps.
Figure 11: Archive of PBS with the number of obstacles and $\lambda_{2}$ as diversity measures.
Number of obstacles vs. $\boldsymbol{\lambda_{2}}$. We run the one-algorithm experiment of PBS with the number of obstacles and $\lambda_{2}$ as the diversity measures to explore the relationship between the two measures and their correlation with the hardness of the generated maps. From Figure 11, we observe that most of the generated maps are hard for PBS to solve. In this case, the QD search is close to convergence, revealing that applying each diversity measure can efficiently generate challenging maps for PBS. Therefore, both diversity measures can quantify the general hardness of generated maps. We use the number of obstacles in our experiments due to its computational efficiency and intuitive interpretation.
A.2 Spatial Arrangement
Figure 12: Archive of PBS with map entropy and KL divergence of the tile pattern distribution as diversity measures.
KL Divergence of Tile Pattern Distribution vs. Map Entropy. We run a one-algorithm experiment with PBS. We use the same objective in Section 4 and use map entropy and KL divergence of the tile pattern distribution as the diversity measures. Figure 12 shows the resulting archive of maps. We observe that, with map entropy less than 0.5, the generated maps are usually composed of a large chunk of empty spaces with a large component of obstacles on the boundary, revealing a simple structure. Compared to maps generated by the KL divergence of the tile distribution in Figure 5 of Section 4, maps generated by map entropy, such as Map (a) and (b), have less diverse patterns. Therefore, we use the KL divergence of tile pattern distribution as the diversity measure.
Figure 13: Archive of PBS with KL divergence of the tile pattern distribution and KL divergence of WL graph feature as diversity measure.
KL Divergence of Tile Pattern Distribution vs. KL Divergence of WL Graph Feature. We run the one-algorithm experiment of PBS with diversity measures of the KL divergence of the tile pattern distribution and the KL divergence of the WL graph feature to explore their correlation and the spatial arrangements in generated maps. From Figure 13 , we observe that all generated maps are well-distributed across the archive, revealing a similar distribution of generated maps with different pairs of KL divergence of the tile pattern distribution and the KL divergence of WL graph feature. Therefore, these two diversity measures are similar. Both quantify the similarity between the generated maps and the pre-defined set of maze maps Stern et al. [2019]. We use the KL divergence of the tile pattern distribution in our experiments since the WL graph feature needs a graph transformation and several iterations of hash transformation, which is more computationally expensive.
Appendix B Implementation and Compute Resource
@@
Implementation. We implement CMA-MAE with Pyribs Tjanaka et al. [2023]. For MAPF algorithms, we use the open source implementations of CBS and EECBS [^3] https://github.com/Jiaoyang-Li/EECBS, PBS [^4] https://github.com/Jiaoyang-Li/PBS, PIBT [^5] https://github.com/Kei18/pibt2, and LTF [^6] https://github.com/AIRI-Institute/learn-to-follow.
Compute Resources. We run our experiments on three machines: (1) a local machine equipped with a 64-core AMD Ryzen Threadripper 3990X CPU and 192 GB of RAM, (2) a local machine with a 64-core AMD Ryzen Threadripper 7980X CPU and 128 GB of RAM, and (3) a high-performance cluster featuring multiple 64-core AMD EPYC 7742 CPUs, each with 256 GB of RAM. All CPU runtimes are measured on machine (1).
Appendix C Maps with similar patterns
Figure 14: Similar maps with evenly distributed short corridor components or large chunks of empty space on which CBS can solve. Maps (a) and (b) are maps with evenly distributed short corridor components. Map (c) and Map (d) are maps with large chunks of empty space on the right, but also long corridors on the left.

### v5

@@ 3.2. Objective
In all experiments, our objective is a function $f:\boldsymbol{X}\rightarrow\mathbb{R}$, where $\boldsymbol{X}$ is the space of all maps. The function $f$ runs one or two MAPF algorithms on $N_{e}$ different MAPF instances with $N_{a}$ agents. For CBS, EECBS, LaCAM3, and PBS, we set a time limit $T$ for each run. For PIBT and LTF, we set a maximum makespan of $M$ timesteps.
@@
One-Algorithm Experiments. We intend to generate maps that are challenging for a given MAPF algorithm $\phi\in A$, where $A=\{\text{CBS},\ \text{EECBS},\ \text{LaCAM3},\ \text{PBS},\ \text{PIBT},\ \text{LTF}\}$ in our experiments. The objective quantifies the empirical hardness of a map $\boldsymbol{x}\in\boldsymbol{X}$ to $\phi$.
@@
For CBS, EECBS, and PBS, the objective is to maximize their CPU runtime because they are exponential-time algorithms, the major limitation being frequently running out of time. We do not consider the solution quality as the objective because CBS and EECBS are proven to be optimal and bounded-suboptimal algorithms. PBS, while being unbounded suboptimal, empirically finds near-optimal solutions. For example, MaAAAI19 shows that PBS never finds solutions over 4% worse than optimal on game maps from the MAPF benchmark. We include LaCAM3 because it can solve almost all instances in the MAPF benchmark. Our objective is to identify more challenging maps for it. Therefore, we maximize CPU runtime to find failure cases for CBS, EECBS, PBS, and LaCAM3. Concretely, suppose that the function $t_{\phi}:\boldsymbol{X}\rightarrow\mathbb{R}_{>0}$ computes the average CPU runtime by running $\phi\in A$ on $N_{e}$ instances, the objective $f$ for the algorithm $\phi$ is $f(\boldsymbol{x})=t_{\phi}(\boldsymbol{x}),\phi\in\{\text{CBS},\ \text{EECBS},\ \text{PBS},\ \text{LaCAM3}\}$. If $\phi$ cannot find solutions within the time limit $T$, we set the CPU runtime to $T$.
@@
$SR_{\phi}:\boldsymbol{X}\rightarrow\mathbb{R}_{>0}$ computes the success rate of running algorithm $\phi$ on one MAPF instance, which is computed as the percentage of agents that successfully reach their goals within the given makespan $M$. $SoC_{\phi}:\boldsymbol{X}\rightarrow\mathbb{R}_{>0}$ computes the sum-of-cost of the solution if $SR=1$. $C$ is a large constant making sure that the RSR score of maps with $SR=1$ always dominates that of maps with $SR<1$. Intuitively, if not all agents reach their goals ($SR<1$), we maximize the success rate. If all agents reach their goals ($SR=1$), we regularize the success rate with the sum-of-cost to quantify the quality of the solution. Let the function $r_{\phi}:\boldsymbol{X}\rightarrow\mathbb{R}_{>0}$ return the average regularized success rate by running algorithm $\phi$ in $N_{e}$ MAPF instances. Concretely, $r_{\phi}(\boldsymbol{x})=\sum_{i=1}^{N_{e}}RSR_{\phi}^{(i)}(\boldsymbol{x})$, where $i$ denotes the $i$-th run of algorithm $\phi$. Then the objective of PIBT and LTF are $f(\boldsymbol{x})=r_{\text{PIBT}}(\boldsymbol{x})$ and $f(\boldsymbol{x})=r_{\text{LTF}}(\boldsymbol{x})$, respectively. We do not consider runtime as the objective of PIBT and LTF because they are both one-step planners and thus their runtime is strongly correlated with success rates and solution quality.
@@
Two-Algorithm Experiments. To compare two given MAPF algorithms, we aim to generate maps that maximize the performance gap between the them. We consider two pairs of comparisons: (1) EECBS vs. PBS and (2) PIBT vs. LTF. We compare EECBS and PBS because (1) both suboptimal algorithms use a two-level search based on collision avoidance, (2) both find near-optimal solutions empirically, and (3) few prior works have compared them thoroughly. We compare PIBT and LTF because (1) both use a pre-defined policy, either rule-based or learned, to move the agents step by step to their goals while avoiding collisions, (2) both run fast but suffer from deadlocks, and (3) while the LTF work learntofollow2024 has compared them on warehouse maps, showing that LTF outperforms PIBT, we want to find cases where PIBT outperforms LTF. To compare EECBS and PBS, we set the objective as the absolute difference in the average CPU runtime. Concretely, the objective of a generated map $\boldsymbol{x}\in\boldsymbol{X}$ is computed as $f(\boldsymbol{x})=|t_{\text{EECBS}}(\boldsymbol{x})-t_{\text{PBS}}(\boldsymbol{x})|$. Similarly, to compare PIBT and LTF, we set the objective as the absolute difference in the regularized success rate, i.e., $f(\boldsymbol{x})=|r_{\text{PIBT}}(\boldsymbol{x})-r_{\text{LTF}}(\boldsymbol{x})|$.
3.3. Diversity Measures
@@
We select diversity measures such that the generated maps are diverse in (1) general hardness towards all algorithms to analyze whether the general hardness of a map aligns with the hardness to a specific algorithm, and (2) spatial arrangement of the obstacles to generate maps of different patterns.
@@
General Hardness. To measure the general hardness of a map, the most commonly used metric is the number of obstacles. The number of obstacles in a map affects the space available to resolve collisions as well as the search space of the algorithms. Therefore, diversifying it would give a range of difficulties. We also consider two metrics that are proposed specifically to reflect the hardness of a map: the standard deviation of Betweenness Centrality (Std of BC) ewing2022BC and $\lambda_{2}$ ren2024MAPFhard.
The Betweenness Centrality of an empty space is the fraction of all possible shortest paths in the map that pass through it. We use the Std of BC here to measure the congestion in the map. The second smallest eigenvalue (known as $\lambda_{2}$) of the normalized Laplacian matrix of different maps is also claimed to be correlated to the difficulty of the maps ren2024MAPFhard.
Spatial Arrangement. In addition to the hardness of the maps, we want to diversify the spatial arrangement of the maps, making the generated maps stylistically diverse. We first consider the KL divergence of the tile pattern distribution fontaine2020illuminating between the tile pattern distribution of the generated map and a set of maps selected from the MAPF benchmark SternSoCS19.
@@
A tile pattern is one possible arrangement of obstacles and empty spaces in a 3 $\times$ 3 grid. To compute the tile pattern distribution of a map, we count the number of possible tile patterns in the map. We then compute the KL divergence between the tile pattern distribution of the map and a fixed distribution calculated from all maze maps [^1] https://movingai.com/benchmarks/grids.html of the single-agent path finding benchmark sturtevant2012benchmarks. The smaller the KL divergence, the more similar in local patterns our generated map is to the selected set of benchmark maps. We choose maze maps to compare the KL divergence with because they empirically yield the most complex patterns.
In addition, we consider the entropy of the tile pattern distribution ZhangNCA2023. The entropy of the tile pattern distribution is a diversity measure to indicate the level of pattern regularization. The lower the entropy, the more regularized patterns the map possesses.
Finally, we consider the KL divergence of the Weisfeiler-Lehman (WL) graph feature Nino2011WLkernal between the generated map and the maze maps. The WL graph feature indicates the similarity between graphs. Similar to the tile pattern distribution, a smaller KL divergence of the WL graph feature implies more similarities between the local patterns of the generated map and the maze maps.
@@
In our experiments, we choose the number of obstacles and KL divergence of tile pattern distribution as the diversity measures. We empirically observe that, among the measures considered, the number of obstacles is the most correlated with the general hardness of the maps, and the KL divergence of tile pattern distribution yields the most diverse map patterns. We justify our choice with experimental results in Appendix B.
4. One Algorithm Analysis
@@
For one-algorithm experiments, we intend to generate diverse difficult maps for CBS, EECBS, PBS, LaCAM3, PIBT, and LTF.
@@
Figure 2. Examples of one-entry space and one-tile entry.
[…]
[Examples of one-entry space and one-tile entry.]One-entry space and one-tile entry are two major patterns that commonly appear in generated maps.
Figure 3. Archive for CBS with sample maps.
[…]
[Archive of CBS]Archive of CBS shows the distribution of difficult maps QD-MAPPER generates with four representative maps.
4.1. Experiment Setup
@@
Hyperparameters. Defining $w$ as the suboptimality bound away from the optimal solution, we use $w=1.5$ for EECBS. For CBS, we use the implementation of EECBS with $w=1$. We set $T=20$ seconds for CBS, EECBS, LaCAM3, and PBS, $M=1000$ for PIBT, and $M=512$ for LTF. For all algorithms, we fix the map size as 32 $\times$ 32. The range for the number of obstacles is $O_{lb}=307,O_{ub}=717]$, which corresponds to 30% to 70% of the map size. For each evaluation, we run the given MAPF algorithm in $N_{e}=5$ different instances, each with $N_{a}$ agents. For CBS, EECBS, and PBS, $N_{a}=50$, for LaCAM3, $N_{a}=100$, and for PIBT and LTF, $N_{a}=150$. We use more agents for LaCAM3, PIBT, and LTF because they empirically have higher success rates in congested MAPF instances. For all algorithms, we evaluated $N_{eval}=10,000$ maps. We present computational costs and resources in Appendix C.
@@
Implementation. We implement CMA-MAE with Pyribs pyribs. For MAPF algorithms, we use the open source implementations of CBS and EECBS [^2] https://github.com/Jiaoyang-Li/EECBS, PBS [^3] https://github.com/Jiaoyang-Li/PBS, LaCAM3444https://github.com/Kei18/lacam3, PIBT [^5] https://github.com/Kei18/pibt2, and LTF [^6] https://github.com/Cognitive-AI-Systems/learn-to-follow.
4.2. Results
Generated Maps. We first define one-entry space and one-tile entry (shown in Figure 2(b)) as features of the maps. A one-entry space represents the empty space surrounded by three obstacles. A one-tile entry represents a single empty space separating two corridors. In Figures 3, 4 and 5, the yellow and dark blue cells in the archives indicate maps with high and low CPU runtime, respectively. In Figures 7 and 8, the yellow and dark blue cells indicate maps with low and high success rates, respectively. The colored tiles in the maps display 10 out of 150 pairs of start and goal locations. While we are primarily interested in maps that are challenging for the algorithms, we also present easy maps as the by-products of the experiments. We summarize the typical patterns of the maps and show more maps with similar patterns in Appendix D.
Figure 4. Archive for EECBS with sample maps.
[…]
[Archive of EECBS]Archive of EECBS shows the distribution of difficult maps QD-MAPPER generates with four representative maps.
CBS. Figure 3 shows the archive of maps for CBS. It is easy to generate challenging maps for CBS. Map (a), Map (c), and Map (d) show the most typical hard maps for CBS. Map (a) and Map (c) contain many long corridors, and Map (d) contains many one-entry spaces with long corridors.
Some maps that CBS can solve are maps with fewer obstacles and a large chunk of empty spaces. Map (b) is one of these typical maps. Map (b) contains a large chunk of empty spaces and short corridor components in between, providing more space for agents to avoid collisions.
EECBS. Figure 4 shows the archive of maps for EECBS. Compared to the archive of CBS, it is harder to generate challenging maps for EECBS. Map (b) and Map (c) are hard maps with typical patterns for EECBS. Map (b) contains a large chunk of empty space, but with many long corridors and one-tile entries. Map (c) contains many long corridors and one-tile entries. EECBS runs out of time on both maps.